A prankster drops a water balloon off the top of a 1,024-ft.-high building.When will the balloon hit the ground?
Round your answer to one decimal place.
The balloon hits the ground after___ seconds.
1
Expert's answer
2012-06-14T08:07:31-0400
1024 ft = 312.115 m
h = g · t^2 / 2
t = sqrt(2 · h / g)
t = sqrt(2 ·
312.115 / 9.81) = sqrt(63.632) = 7.977 s
The balloon hits the ground after
7.977 (approx. 8.0) seconds.